# Index Terms—  LiDAR Point Clouds, 2D Range-View Projection, Many-to-One Mapping, Semantic Segmentation

### v1

@@ Abstract
Range-view projection provides an efficient method for transforming 3D LiDAR point clouds into 2D range image representations, enabling effective processing with 2D deep learning models. However, a major challenge in this projection is the many-to-one conflict, where multiple 3D points are mapped onto the same pixel in the range image, requiring a selection strategy. Existing approaches typically retain the point with the smallest depth (closest to the LiDAR), disregarding semantic relevance and object structure, which leads to the loss of important contextual information. In this paper, we extend the depth-based selection rule by incorporating contextual information from both instance centers and class labels, introducing two mechanisms: Centerness-Aware Projection (CAP) and Class-Weighted-Aware Projection (CWAP). In CAP, point depths are adjusted according to their distance from the instance center, thereby prioritizing central instance points over noisy boundary and background points. In CWAP, object classes are prioritized through user-defined weights, offering flexibility in the projection strategy. Our evaluations on the SemanticKITTI dataset show that CAP preserves more instance points during projection, achieving up to a 3.1% mIoU improvement compared to the baseline. Furthermore, CWAP enhances the performance of targeted classes while having a negligible impact on the performance of other classeshttps://github.com/s110m/Contextual_Range-View_Projection.
@@ 1 Introduction
Deep learning approaches, including both 3D and 2D models, have emerged as the state-of-the-art for point cloud processing across a wide range of perception tasks. Examples include semantic segmentation [14, 1], which assigns a semantic label to each point; panoptic segmentation [7], which combines semantic labeling with instance identification; and object detection [6], which involves localizing and classifying objects in a scene.
Point-based networks have been developed to directly process raw point clouds, offering a principled solution for unstructured 3D data [15, 16]. However, their high memory and computational costs limit scalability in large-scale or real-time applications. Voxel-based methods discretize space into regular grids to leverage 3D convolutions, improving compatibility with CNNs but suffering from high memory usage and inefficiency caused by data sparsity and the expense of 3D operations [24]. Thus, 3D deep learning methods face significant challenges in meeting the real-time and resource constraints of autonomous driving [20, 21].
@@
Extensive experiments on the SemanticKITTI dataset demonstrate the effectiveness of our approach. CAP preserves more instance points during projection and achieves up to a 3.1% mIoU improvement over baseline methods in semantic segmentation. Moreover, CWAP improves the performance of user-targeted classes through weighted prioritization while maintaining a negligible impact on other classes.
@@ 4.1 Centerness-Aware Projection (CAP)
$\mu_{x}=\tfrac{x_{\min}+x_{\max}}{2},\quad\mu_{y}=\tfrac{y_{\min}+y_{\max}}{2},\quad\mu_{z}=\tfrac{z_{\min}+z_{\max}}{2}\cdot$
@@ 4.2 Class-Weighted-Aware Projection (CWAP)
Table 1: SemanticKITTI comparison between depth-based, CAP, and CWAP range-view projections on the semantic segmentation task. $\text{mIoU}^{\text{Inst}}$ is the mean over classes 1–8 (Instance classes), and $\text{mIoU}^{\text{Stuff}}$ is the mean over classes 9–19 (Stuff classes). We used the RangeViT architecture for all methods.
Projection car bicycle motorcycle truck other-veh. person bicyclist motorcyclist road parking sidewalk other-gr. building fence vegetation trunk terrain pole t-sign mIoU $\text{mIoU}^{\text{Inst}}$ $\text{mIoU}^{\text{Stuff}}$
Depth 93.4 36.1 45.9 56.8 31.7 54.3 65.4 0.0 94.6 42.7 81.9 0.3 87.8 56.1 86.2 64.8 73.2 62.8 45.5 56.8 48 63.3
CAP 93.8 37.0 46.7 69.1 39.3 58.4 64.6 0.1 94.6 48.1 82.1 1.1 86.6 55.3 84.9 63.1 70.2 62.9 45.0 58 51.1 63
CWAP 93.3 35.5 48.5 77.5 33.9 54.6 67.6 0.0 94.7 44.4 81.8 0.8 87.7 57.1 84.9 64.4 70.0 63.4 44.2 58.1 51.4 63
@@ 5.4 Semantic Segmentation
Table 1 reports the impact of our projection methods (CAP and CWAP) on the semantic segmentation task. CAP demonstrates consistent gains across nearly all instance-level classes, achieving an overall improvement of $3.1\%$ mIoU on instance-level categories. The mIoU for stuff classes decreases by $0.3\%$. This drop can be attributed to two factors: (1) CAP assigns scores only to instance objects, since center points can be defined for them, and (2) changes in the IoU of certain categories indirectly affect others, as the evaluation metric is sensitive to the balance between true positives and false positives across classes. For the CWAP strategy, we conduct targeted experiments by assigning a weight of 1-1 to specific object classes (motorcycle, truck, and bicyclist) while setting all others to zero. This configuration yields a notable performance gain for the emphasized classes. For instance, performance on the truck class improves by $8.4\%$ and $20.7\%$ compared to CAP and depth-based range-view projection, respectively. In addition, CWAP improves the mIoU on instance-level categories by $0.3\%$ and $3.4\%$ compared to CAP and depth-based methods. However, with CWAP, the performance of certain instance objects and stuff classes decreases, as the adjustments made to prioritize specific object categories can create ripple effects on other instances and background classes.
@@ 6 Conclusion
Experimental results on SemanticKITTI demonstrate that improving projection quality during training leads to consistent gains at inference, with CAP achieving a 3.1% mIoU improvement on instance-level categories and CWAP providing targeted gains for selected classes.

### v2

@@ Abstract
Range-view projection provides an efficient method for transforming 3D LiDAR point clouds into 2D range image representations, enabling effective processing with 2D deep learning models. However, a major challenge in this projection is the many-to-one conflict, where multiple 3D points are mapped onto the same pixel in the range image, requiring a selection strategy. Existing approaches typically retain the point with the smallest depth (closest to the LiDAR), disregarding semantic relevance and object structure, which leads to the loss of important contextual information. In this paper, we extend the depth-based selection rule by incorporating contextual information from both instance centers and class labels, introducing two mechanisms: Centerness-Aware Projection (CAP) and Class-Weighted-Aware Projection (CWAP). In CAP, point depths are adjusted according to their distance from the instance center, thereby prioritizing central instance points over noisy boundary and background points. In CWAP, object classes are prioritized through user-defined weights, offering flexibility in the projection strategy. Our evaluations on the SemanticKITTI dataset show that CAP preserves more instance points during projection, achieving up to a 4.0 mIoU points improvement over the baseline.
@@ 1 Introduction
Deep learning approaches, including both 3D and 2D models, have emerged as the state of the art for point cloud processing across a wide range of perception tasks. Examples include semantic segmentation [16, 1], which assigns a semantic label to each point; panoptic segmentation [8], which combines semantic labeling with instance identification; and object detection [7], which involves localizing and classifying objects in a scene.
Point-based networks have been developed to directly process raw point clouds, offering a principled solution for unstructured 3D data [17, 18]. However, their high memory and computational costs limit scalability in large-scale or real-time applications. Voxel-based methods discretize space into regular grids to leverage 3D convolutions, improving compatibility with CNNs but suffering from high memory usage and inefficiency caused by data sparsity and the expense of 3D operations [27]. As a result, 3D deep learning methods face significant challenges in meeting the real-time and resource constraints of autonomous driving. [22, 23].
@@
Extensive experiments on the SemanticKITTI dataset demonstrate the effectiveness of our approach. CAP preserves more instance points during projection and achieves up to a 4.0 mIoU point improvement over baseline methods in semantic segmentation. Moreover, CWAP improves the performance of user-targeted classes through weighted prioritization while maintaining a negligible impact on other classes.
@@ 4.1 Centerness-Aware Projection (CAP)
$\mu_{x}=\tfrac{x_{\min}+x_{\max}}{2},\quad\mu_{y}=\tfrac{y_{\min}+y_{\max}}{2},\quad\mu_{z}=\tfrac{z_{\min}+z_{\max}}{2}.$
@@ 4.2 Class-Weighted-Aware Projection (CWAP)
Table 1: SemanticKITTI comparison between depth-based, CAP, and CWAP range-view projections methods on the semantic segmentation task. $\text{mIoU}^{\text{Inst}}$ is the mean over classes 1–8 (Instance classes), and $\text{mIoU}^{\text{Stuff}}$ is the mean over classes 9–19 (Stuff classes). We used the RangeViT architecture for all methods.
Method car bicycle motorcycle truck other-veh. person bicyclist motorcyclist road parking sidewalk other-gr. building fence vegetation trunk terrain pole t-sign mIoU $\text{mIoU}^{\text{Inst}}$ $\text{mIoU}^{\text{Stuff}}$
RangeNet++ [16] 91.4 25.7 34.4 25.7 23.0 38.3 38.8 4.8 91.8 65.0 75.2 27.8 87.4 58.6 80.5 55.1 64.6 47.9 55.9 52.2 35.3 64.5
SqueezeSegV3 [24] 92.5 38.7 36.5 29.6 33.0 45.6 46.2 20.1 91.7 63.4 74.8 26.4 89.0 59.4 82.0 58.7 65.4 49.6 58.9 55.9 42.8 61.7
SalsaNext [5] 91.9 48.3 38.6 38.9 31.9 60.2 59.0 19.4 91.7 63.7 75.8 29.1 90.2 64.2 81.8 63.6 66.5 54.3 62.1 59.5 48.5 67.5
KPRNet [9] 95.5 54.1 47.9 23.6 42.6 65.9 65.0 16.5 93.2 73.9 80.6 30.2 91.7 68.4 85.7 69.8 71.2 58.7 64.1 63.1 51.4 71.6
RangeViT [1] 95.4 55.8 43.5 29.8 42.1 63.9 58.2 38.1 93.1 70.3 80.0 32.5 92.0 69.0 85.3 70.6 71.2 60.8 64.7 64.0 53.4 71.8
CAP 95.8 56.2 46.7 49.1 44.2 64.1 64.6 38.2 94.6 70.2 82.1 31.2 91.8 68.5 84.9 70.3 70.2 60.2 64.0 65.8 57.4 71.6
CWAP 93.3 35.5 48.5 58.4 40.9 62.6 55.9 38.0 94.7 70.1 81.8 30.1 90.7 68.2 84.9 70.4 70.0 61.2 64.2 64.1 54.1 71.5
@@ 5.4 Semantic Segmentation
Table 1 reports the impact of our projection methods (CAP and CWAP) on the semantic segmentation task. CAP demonstrates consistent gains across nearly all instance-level classes, achieving an improvement of 4.0 mIoU points on instance-level categories. The mIoU for stuff classes decreases by 0.2 mIoU points. This drop can be attributed to two factors: (1) CAP assigns scores only to instance objects, since center points can be defined for them, and (2) changes in the IoU of certain categories indirectly affect others, as the evaluation metric is sensitive to the balance between true positives and false positives across classes. For the CWAP strategy, we conduct targeted experiments by assigning a weight of 1-1 to specific object classes (motorcycle, truck, and bicyclist) while setting all others to zero. This configuration yields a notable performance gain for the emphasized classes. For instance, performance on the truck class improves by 9.3 mIoU points and 28.6 mIoU points compared to CAP and depth-based range-view projection in RangeViT, respectively. In addition, CWAP improves the mIoU on instance-level categories by 0.7 mIoU points compared to depth-based RangeViT method. However, with CWAP, the performance of certain instance objects and stuff classes decreases, as the adjustments made to prioritize specific object categories can create ripple effects on other instances and background classes.
@@ 6 Conclusion
Experimental results on SemanticKITTI demonstrate that improving projection quality during training leads to consistent gains at inference, with CAP achieving a 4.0 mIoU point improvement on instance-level categories and CWAP providing targeted gains for selected classes.